Define: R^3→R^3 by


T(x,y,z)=(x-y+z,x+y,y+z)


Let v1= (1,1,1), v2= (0,1,1), v3= (0,0,1). Find a matrix of T with respect to the basis {v1,v2,v3}. Futher check T is invertible or not.


1
Expert's answer
2021-06-02T13:55:57-0400

"V_1=[1,1,1]"

"V_2=[0,1,1]"

"V_3=[0,0,1]"


"T=[x-y+z,x+y,y+z]"


According to this the new


"V_1'=[1,2,2]"

"V_2'=[0,1,2]"

"V_3'=[1,0,1]"


So the matrix T form is as follows:


"T=" "\begin{bmatrix}
1 & 2 & 2 \\
0 & 1 & 2\\
1 & 0 & 1
\end{bmatrix}"


For the invertible matrix the determinant of the matrix is = 0


Det(T) = "1(1-0)-2(0-2)+2(0-1)=1+4-2=3"


As the determinant of the matrix is not equal to zero hence the given matrix is invertible.